OpenStudy (joshoyen):

This is confusing. Please help. List the zeros, degrees, and end behavior of each functions. Then draw a general sketch of the function. 1. y = (x+3)(x-2) 2. y = -(x+2)(x-4)

OpenStudy (butterflydreamer):

To find the zeros of a function, you set y = 0 and solve for x :) Degree of a function is the largest exponent (power) of the variable... So for e.g. \[y = x^4 + 2x - 7\] the degree would be 4.

OpenStudy (joshoyen):

So for #1, the degree would be 2?

OpenStudy (butterflydreamer):

yes! :)

OpenStudy (joshoyen):

Oh, okay. Thank you!

OpenStudy (butterflydreamer):

no problemo :D

OpenStudy (butterflydreamer):

Do you know how to find the end behavior and sketch ?

OpenStudy (joshoyen):

Uh, lol not really sure

OpenStudy (butterflydreamer):

okay, so the end behaviour is basically just asking which way the ends of the function is facing...So i'll use an example. If we had the function y = ( x - 4 ) ( x + 2 ) and they wanted me to list the zeros, degrees and end behaviour and sketch i would firstly do this: Set y = 0 to find the ZEROS 0 = ( x - 4 ) ( x + 2 ) This means that x - 4 = 0 or x + 2 = 0 Therefore x = 4 or -2 right? Next, just by looking at the function.. you should be able to tell that it is a quadratic function (so the highest degree would be 2 bc x^2). Then if we look at the function y = (x - 4) (x + 2), if i EXPAND this it will become: y = x^2 + 2x - 4x - 8 y = x^2 - 2x - 8 and since x^2 is POSITIVE (so it isn't -x^2), i know this means that the sketch will be a "happy face sketch" meaning that the ends will both be increasing:  See how when x = -2 , y = 0 and when x = 4, y = 0 (you can think of these as your x-intercepts) and these points will show you where the graph will cut the x-axis. If you look at the arrow heads (ends of the graph) can you see how both are going upwards? this is because of the POSITIVE x^2 :) hope this helps